The daily demand for a local newspaper at a street vendor has the following probabilities:

Demand

0     1

2     3     4     5     6

Probability

.1    .2

.3    .1    .1    .1    .1

The news vendor must pay 50 cents for each copy. Each newspaper sells for $1 and unsold copies are returned to the distributor for a 5 cent refund.

a.[7 marks] When 6 copies are stocked, calculate the profit (or loss) for each possible level of demand.

b.[2 marks] When 6 copies are stocked, what is the expected profit?

c.[6 marks] How many copies should the news vendor stock to maximize expected profit?
Answer #1

a.  

6 copies are stocked.

At demand level 0, loss = 6\times50 cents - 6\times5 cents = 300 cents - 30 cents = 270 cents = $2.70

At demand level 1, loss = 6\times50 cents - $1 - 5\times5 cents = 300 cents -100 cents - 25 cents = 175 cents = $1.75

At demand level 2, loss = 6\times50 cents - 2\times$1 - 4\times5 cents = 300 cents -200 cents - 20 cents = 80 cents = $0.80

At demand level 3, profit = 3\times$1 + 3\times5 cents - 6\times50 cents = 300 cents +15 cents - 300 cents = 15 cents = $0.15

At demand level 4, profit = 4\times$1 + 2\times5 cents - 6\times50 cents = 400 cents + 10 cents - 300 cents = 110 cents = $1.10

At demand level 5, profit = 5\times$1 + 5 cents - 6\times50 cents = 500 cents + 5 cents - 300 cents = 215 cents = $2.15

At demand level 6, profit = 6\times$1 - 6\times50 cents = 600 cents - 300 cents = 300 cents = $3

.

b.

Demand 0 1 2 3 4 5 6
Profit -$2.70 -$1.75 -$0.80 $0.15 $1.10 $2.15 $3
Probability 0.1 0.2 0.3 0.1 0.1 0.1 0.1

Expected profit = (-$2.70)\times0.1 + (-$1.75)\times0.2 + (-$0.80)\times0.3 + $0.15\times0.1 + $1.10\times0.1 + $2.15\times0.1 + $3\times0.1 = -$0.22

Thus the expected loss is 22 cents.

.

c.

Let the news vendor stock n copies, n=1,2,3,4,5,6

When n=1 :-

At demand level 0, profit = 5 cents - 50 cents = -45 cents

At demand level 1, profit = $1 - 50 cents = 50 cents

Expected profit = -45\times0.1 cents + 50\times0.2 cents = 5.5 cents

When n=2 :-

At demand level 0, profit = 2\times5 cents - 2\times50 cents = -90 cents

At demand level 1, profit = $1 + 5 cents - 2\times50 cents = 5 cents

At demand level 2, profit = 2\times$1 - 2\times50 cents = 100 cents

Expected profit = -90\times0.1 cents + 5\times0.2 cents + 100\times0.3 cents = 38 cents

When n=3 :-

At demand level 0, profit = 3\times5 cents - 3\times50 cents = -135 cents

At demand level 1, profit = $1 + 2\times5 cents - 3\times50 cents = -40 cents

At demand level 2, profit = 2\times$1 + 5 cents - 3\times50 cents = 45 cents

At demand level 3, profit = 3\times$1 - 3\times50 cents = 150 cents

Expected profit (in cents) = -0.1\times135 cents - 0.2\times40 cents + 0.3\times45 cents + 0.1\times150 cents = 7 cents

When n=4 :-

At demand level 0, profit = 4\times5 cents - 4\times50 cents = -180 cents

At demand level 1, profit = $1 + 3\times5 cents - 4\times50 cents = -85 cents

At demand level 2, profit = 2\times$1 + 2\times5 cents - 4\times50 cents = 10 cents

At demand level 3, profit = 3\times$1 + 5 cents - 4\times50 cents = 105 cents

At demand level 4, profit = 4\times$1 - 4\times50 cents = 200 cents

Expected profit (in cents) = -0.1\times180 cents - 0.2\times85 cents + 0.3\times10 cents + 0.1\times105 cents + 0.1\times200 cents = -1.5 cents

When n=5 :-

At demand level 0, profit = 5\times5 cents - 5\times50 cents = -225 cents

At demand level 1, profit = $1 + 4\times5 cents - 5\times50 cents = -130 cents

At demand level 2, profit = 2\times$1 + 3\times5 cents - 5\times50 cents = -35 cents

At demand level 3, profit = 3\times$1 + 2\times5 cents - 5\times50 cents = 60 cents

At demand level 4, profit = 4\times$1 + 5 cents - 5\times50 cents = 155 cents

At demand level 5, profit = 5\times$1 - 5\times50 cents = 250 cents

Expected profit ( in cents) = -0.1\times225 cents - 0.2\times130 cents - 0.3\times35 cents + 0.1\times60 cents + 0.1\times155 cents + 0.1\times250 cents = -12.5 cents

When n=6 , expected profit = -$0.22 = -22 cents

Thus it is seen that expected profit is maximum when n=2

The news vendor should stock 2 copies to maximize expected profit.